roger studied for x hours. her sister studied five more than roger. between both of them, they studied 30 hours. how much did each study?
Found 2 solutions by mananth, texttutoring:
Answer by mananth(16949)

roger = xhours
sister = x+5 hours
..
x+x+5 =30
2x+5=30
-5 -5
2x=25
/2
x= 12.5 hours. roger time
sister = 17.5 hours
...
m.ananth@hotmail.ca

Answer by texttutoring(324)

Roger studied for x hours
Her sister studied for (x+5) hours.

Together, they studied for 30 hours in total. You need to set up an algebra equation to model this question:

x + (x+5) = 30

Now collect like terms:

2x + 5 = 30
2x = 30 -5
2x = 25
x = 12.5

So Roger studied for 12.5 hours. That means his sister must have studied for 12.5+5 = 17.5 hours.

Check your work. Does 12.5+17.5 = 30?

Yes, it does! So that means this must be correct.